OpenStudy (anonymous):

"What is the ninth root of 32^3 in simplified radical form with the smallest possible index?" A. 2 * square root of 4 B. 2 * third root of 2 C. 2 * third root of 4 D. 4 * third root of 3

zepdrix (zepdrix):

We can write the 9th root like this, as a rational expression. \[\large \left(32^3\right)^{1/9}\] Using our laws of exponents, we multiply the powers,\[\large 32^{3/9}\]Which simplifies to,\[\large 32^{1/3}\] To simplify further we want to find a factor of 32 that is a `perfect cube`.

zepdrix (zepdrix):

\[\large 32 = 8\cdot 4\]Hmm it has a factor of 8. \[\large 8=2\cdot2\cdot2=2^3\] 8 is a perfect cube! We can take the cube root of 8.\[\large (32)^{1/3} \qquad = \qquad (8\cdot4)^{1/3} \qquad = \qquad 2(4)^{1/3}\]Understand how that works? :)

OpenStudy (anonymous):

Thank you so much! (: